OpenStudy (anonymous):

Find the eccentricity, directrices and focal width of the ellipse 2x^2=1-y^2

OpenStudy (anonymous):

ok so 2x^2+y^2=1

OpenStudy (anonymous):

what to do next?

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

i was trying to get it in the from x^2/a^2+y^2/b^2=1

OpenStudy (mathmate):

To find a and b: compare \(x^2/a^2+y^2/b^2=1\) with \(x^2/(1/\sqrt 2)^2+y^2=1\)

OpenStudy (anonymous):

ok y couldnt we just put x^2/(1/2)+y^2=1 or thats not correct?

OpenStudy (mathmate):

we have a^2=(1/2) so a=sqrt(1/2)=1/sqrt(2)=sqrt(2)/2

OpenStudy (mathmate):

@blow_pop2000 Do you mind finding someone to chat with in the chat area, like what you're doing in the History section? Thank you.

OpenStudy (mathmate):



OpenStudy (mathmate):

we have a=sqrt(2)/2, b=1

OpenStudy (anonymous):

ok so a=sq.rt 2/2 then a^2=2/4 or 1/2 ok

OpenStudy (mathmate):

exactly, that's how you get back 2x^2+y^2=1

OpenStudy (anonymous):

so eccentricity would be c/b where c^2= b^2-a^2 c^2=1-1/2 c^2=1/2 c=sq.rt2/2 so eccentricity =(sq.rt 2/2)/1

OpenStudy (anonymous):

right?

OpenStudy (mathmate):

exactly! e=sqrt(2)/2

OpenStudy (mathmate):

and c=sqrt(2)/2 as well.

OpenStudy (mathmate):

So now you'll need the directrices.